OpenStudy (anonymous):

Is my answer right? :) Solve. 2x2 + 3x = -8 -3 _+_ 11i/4

OpenStudy (sirm3d):

here's the correct answer \[\large \frac{ -3 \pm \sqrt{55}i }{ 4 }\] here's the part where you made the mistake:\[\large \frac{ -3 \pm \sqrt{9-64} }{ 4 }=\frac{ -3 \pm \color{red}{(i \sqrt9 +i \sqrt{64})} }{ 4 }=\frac{ -3 \pm \color{red}{11i} }{ 4 }\]

OpenStudy (anonymous):

You're so kind, wow thank you!!!!